# EEG2GAIT: A Hierarchical Graph Convolutional Network for EEG-based Gait Decoding

### v1

@@ 3.4 Hierarchical GCN Pyramid
The HGP learns embedding $g^{i}$ for each subgraph $\hat{{X_{i}}}\in\mathbb{R}^{C\times F}$ to capture spatial features of the signal. A base graph encoder $\Phi_{g}(\cdot)$ is utilized to learn graph representations. In this paper, we use GCN [kipf2022semi] as the base graph encoder:
$\Phi_{g}(\mathbf{H}^{(l+1)},\mathbf{A})=\sigma\left(\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-\frac{1}{2}}\mathbf{H}^{(l)}\mathbf{W}^{(l)}+\mathbf{b}^{(l)}\right),$
@@
  • $\tilde{\mathbf{A}}=\mathbf{A}+\mathbf{I}$ is the adjacency matrix with added self-loops;
  • $\tilde{\mathbf{D}}$ is the degree matrix of $\tilde{\mathbf{A}}$, i.e., $\tilde{D}_{ii}=\sum_{j}\tilde{A}_{ij}$;
@@
In practice, $\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-\frac{1}{2}}$ represents the normalized adjacency matrix, which is used to preserve the scale of feature representations after each graph convolution.
To modulate the diversity of brain region connections for different cognitive processes, we propose to use multiple hierarchical GCNs, i.e., $\{\Phi^{0}_{g}(\cdot),\Phi^{1}_{g}(\cdot),\dots,\Phi^{i}_{g}(\cdot)\}$, with learnable adjacency matrices $\{A^{0},A^{1},\dots,A^{i}\}$. Each $A^{i}$ learns the graph connections corresponding to a particular cognitive process. For each $\Phi^{i}_{g}(\cdot)$, stacking different GCN layers allows learning different levels of node cluster similarity. Intuitively, for localized connections, such as electrodes within the same brain functional area, a deeper GCN can achieve consistent representations among these nodes. However, for global connections across different brain regions, a shallower GCN can aggregate information between these areas without oversmoothing. The deeper $\Phi^{i}_{g}(\cdot)$ becomes, the higher the level in the feature pyramid its output represents.
@@ 3.5 Global Spatial Learner
The GSL processes these combined features using a depth-wise grouped convolution[krizhevsky2012imagenet] with a kernel size of $(C,1)$. This operation efficiently learns spatial filters specific to each temporal filter, enabling the extraction of frequency-specific spatial patterns. To further enhance the feature representations, batch normalization[ioffe2015batch] is applied along the feature map dimension, followed by an exponential linear unit (ELU)[clevert2015fast] activation function. Dropout[srivastava2014dropout] with a probability of 0.5 is used to prevent overfitting when training on small datasets. Finally, an average pooling layer of size $(1,3)$ reduces the temporal resolution. As a result, the output size of the GSL is $(25,1,T//3)$.
@@ 3.6 Feature Fusion Layers
After the three convolutional blocks, the output passes through a dropout layer with a rate of 0.5 to mitigate overfitting. The final output of the Feature Fusion Layers has a size of $(200,1,T//81)$, ready for subsequent temporal modeling.
@@ 3.7 Global Temporal Learner
The Global Temporal Learner (GTL) captures temporal dynamics using a Multi-head Self-Attention (MSA) mechanism[vaswani2017attention], which preserves the input-output shape consistency for efficient processing. Residual connections are employed to maintain the integrity of the input data and ensure stable gradient flow during training.
[…]
@@
$X_{\text{attention}}=\text{linear}(\text{Concat}(\text{head}_{1},\ldots,\text{head}_{h}))\quad(5)$
Here, $N_{\text{head}}$ is the number of attention heads, $d_{K}$ is the output dimension of each attention head, and $d_{\text{model}}$ is the final feature dimension after projection.
@@ 3.9.1 Time-Domain Loss
$L_{\text{time}}=\frac{1}{n}\sum_{i=1}^{n}\left(\hat{y}_{i}-y_{i}\right)^{2}$
@@
$L_{\text{time\_reward}}=L_{\text{time}}+\beta\cdot\log\left(1-e^{-L_{\text{time}}}+\epsilon\right)$
@@ 3.9.2 Frequency-Domain Loss
$L_{\text{freq}}=\frac{1}{n}\sum_{i=1}^{n}\left|\hat{Y}_{i}-Y_{i}\right|$
where $\hat{Y}(f)=\text{DFT}(\hat{y}(t))$ and $Y(f)=\text{DFT}(y(t))$ are the predicted and true signals in the frequency domain. Similar to the time-domain loss, an Encouraging Loss term is applied:
$L_{\text{freq\_reward}}=L_{\text{freq}}+\beta\cdot\log\left(1-e^{-L_{\text{freq}}}+\epsilon\right)$
@@ 3.9.3 HTSR Loss
$L_{\text{total}}=\alpha\cdot L_{\text{freq\_reward}}+(1-\alpha)\cdot L_{\text{time\_reward}}$
@@ 4.1.1 Gait-EEG Dataset
To investigate the brain mechanisms involved in walking, we collected a new dataset, Gait-EEG Dataset (GED), recording brain activity along with simultaneous lower-limb joint angles natural walk on level ground. The dataset contains the recordings from 50 able-bodied participants (25 males, 25 females; aged 21 to 46, mean age 28.4, standard deviation 5.2), with no history of neurological disorders or lower limb pathologies. Participants engaged in two independent level-ground walking experiment sessions, with every session comprising three identical walking blocks. Each block included approximately 40 trials, with each trial representing EEG signals and synchronized lower-limb joint angles as the participant walked straight for 7.7 meters. The experiment protocol for each block is shown in Fig. 2. Sessions were spaced at least three days apart. The dataset includes synchronized recordings from a 60-channel active EEG, a 4-channel electrooculogram (EOG), along with measurements from six joint angle sensors (bilateral hips, knees, and ankles).
This study has been reviewed and approved by the Institutional Review Board (IRB-2021-709) of Nanyang Technological University, ensuring compliance with applicable legislation, ethical and safety requirements in Singapore. All participants have provided informed consents before the experiment.
Figure 2: Experiment protocol for one block.
@@ 4.1.2 Open-access Dataset
To further validate our proposed method, we also conducted experiments using an open-access dataset, Mobile Brain-body imaging(MoBI) dataset[he2018mobile]. The MoBI dataset includes recordings from eight participants across three sessions each. For each session, participants walked at a constant speed on a treadmill for a total of 20 minutes. During the first 15 minutes, participants received real-time feedback of their gait via an on-screen avatar mirroring their actual movements. In the final 5 minutes, the avatar’s gait feedback was driven by decoded EEG signals.
@@ 4.2 Data Preprocessing
Prior to analysis, the EEG data from both datasets underwent the same preprocessing steps. Initially, a minimum-phase filter was employed for band-pass filtering between 0.1 and 48 Hz. Subsequently, a common average reference was applied. The data were then downsampled from 100 Hz, followed by Independent Component Analysis (ICA). All components obtained from ICA were subjected to the computation of five metrics[mognon2011adjust]: Spatial Average Difference, Spatial Eye Difference, Generic Discontinuities Spatial Feature, Maximum Epoch Variance, and Temporal Kurtosis.
A threshold for each metric was determined through maximum likelihood estimation. If any of the five metrics for a given component exceeded three times the threshold, that component was discarded. On average, 8.96 components were removed for each session in GED and 9.54 components were removed for each session in MoBI. Subsequently, the data from the EOG channels were removed, resulting in a final EEG dataset consisting of 59 channels for GED and MoBI dataset.
Finally, a local Laplacian filtering was applied to enhance local activity within a specified area. It achieves this by computing the differences between the values of the target channel and the average of values from its neighbor channels. For each EEG channel, neighboring channels within a 30 mm radius are retained as its neighbors. This method effectively emphasizes the local features and diminishes the interference from distant sources. The mathematical representation of the local Laplacian filtering process is given by the following formula:
@@
where:
  • $V^{\prime}_{k}$ is the value of the $k^{th}$ channel after applying the Laplacian filter.
  • $V_{k}$ is the original value at the $k^{th}$ channel.
  • $N$ is the number of neighbors used for calculating the average.
  • $V_{l}$ represents the original values of the surrounding channels.
@@ 4.3.1 Gait-EEG Dataset
The last 20 trials of the third block were divided, with the first 5 trials serving as the validation set and the subsequent 15 trials being concatenated along the time dimension to build the test set, while the remaining trials formed the training set. Since EEG2GAIT is a session-specific model, we separately trained EEG2GAIT models using training sets from each session and tested them with the test sets from their respective sessions.
@@ 4.3.2 MoBI Dataset
For each session, we use the first 15 minutes of walking data without closed-loop BCI control as the training set and the last 5 minutes of walking data with closed-loop BCI control as the test set. Additionally, the last 10% of the training set, which is the final 1.5 minutes of data, is set aside as the validation set.
@@ 4.4 Evaluation Metric
We evaluated the efficacy of EEG2GAIT by comparing the predicted angles of six joints with their actual recorded angles, using their correlation ($r$ value) and coefficient of determination ($R^{2}$ score) as the evaluation metrics:
@@
where $y$ denotes the actual joint angle, and $\hat{y}$ represents the predicted angle. The covariance between any two variables A and B is expressed as $cov(A,B)$, while $\sigma(A)$ signifies the standard deviation of A. Each sequence represents the data collected from a single trial, capturing the observations over a specific period. In this case, the sequence contains $n$ data points, which corresponds to $\frac{n}{100}$ seconds of data, given that the data was sampled at a frequency of 100 Hz. $\hat{y}_{i}$ refers to the value at the $i^{th}$ position. Lastly, $\bar{y}$ stands for the average of the actual joint angles. The $r$ value reflects the consistency in trend between the predicted and actual gait. The $R^{2}$ score reflects the second-order absolute error between the predicted and actual gait, providing insight into the predictive performance of the model.
@@ 4.5 Implementation and Hyperparameter Settings
5 Results and Analysis
In this section, we validate the performance of the proposed method on the Gait-EEG dataset and MoBI dataset[he2018mobile] and compared with several state-of-the-art deep learning and machine learning algorithms in the brain-computer interface (BCI) domain, including ContraWR[yang2021self], FFCL[li2022motor], TSception[ding2022tsception], Temporal Convolutional Network (TCN)[ingolfsson2020eeg], ST-Transformer[song2021transformer], EEGConformer[song2023eeg], SPaRCNet[jing2023development], EEGNet[lawhern2018eegnet], and deepConvNet[schirrmeister2017deep]. An ablation study was conducted to demonstrate the contributions of the HGP module and HTSR loss. Saliency mapping across EEG channels identified specific channels with notably higher contributions to the decoding task, offering a spatial interpretation of the neural correlates involved in lower-limb motor decoding as well as validating the obtained features. This confirms that the model effectively leverages signals from motor-related areas. Additionally, we illustrate the overlay of decoding results across different models for the gait cycle, offering a more intuitive display of EEG2GAIT’s superior performance.
Fig. 3 illustrates the visualization of the predicted and actual knee joint angles for the left knee during the test set of participant 2’s second session ($r$ = $0.962$, $R^{2}$ = $0.918$). The gait cycles were segmented using peak detection on the goniometer-measured angles, with each peak representing a boundary between consecutive gait cycles. The segment between two consecutive peaks was defined as a single gait cycle, which was subsequently interpolated to a fixed length of 400 time points for standardization.
Figure 3: Visualization of the actual and predicted left knee joint angles for the test set in participant 2’s second session. The goniometer-measured angles were segmented into gait cycles using peak detection, with each gait cycle interpolated to 400 time points. The blue solid line represents the actual mean angle, and the blue shaded area indicates the standard deviation of the measured angles across all gait cycles. The red dashed line denotes the predicted mean angle, while the red shaded area shows the standard deviation of the predictions.
5.1 Model Performance
The results of all metrics on the test sets are presented to evaluate the performance of EEG2GAIT in comparison to baseline methods on the MoBI and GED dataset. EEG2GAIT achieved an $r$ value of 0.935 ($R^{2}$ = 0.865) on the GED dataset with 50 subjects and an $r$ value of 0.713 ($R^{2}$ = 0.484) on the MoBI dataset with 8 subjects, outperforming the baseline methods. All baseline methods were trained according to the strategy described in Section 4.3. A summary of the results for EEG2GAIT and the baseline methods on the different datasets is provided in Table 1 and 2. Overall, EEG2GAIT demonstrated the best performance across all metrics among all methods, with a lower standard deviation than most methods.
Table 1: Prediction performance of different models on the MoBI dataset, indicated by mean $\pm$ standard deviation of the correlations ($r$) and coefficients of determination ($R^{2}$) between predictions and true signals, averaged over the six joints. The symbol $\pm$ indicates the mean value followed by its standard deviation. The highest mean value in each column is marked in bold, while the lowest standard deviation in each column has been underlined.
Model Name $r$ value $R^{2}$ score
ContraWR[yang2021self] 0.314$\pm$0.1693 0.0427$\pm$0.1117
FFCL[li2022motor] 0.4407$\pm$0.1996 0.1413$\pm$0.2285
TSception[ding2022tsception] 0.5527$\pm$0.1982 0.1918$\pm$0.3661
TCN[ingolfsson2020eeg] 0.6099$\pm$0.1658 0.3423$\pm$0.2396
ST-Transformer[song2021transformer] 0.676$\pm$0.1996 0.4099$\pm$0.3183
EEGConformer[song2023eeg] 0.6027$\pm$0.1948 0.3325$\pm$0.2816
SPaRCNet[jing2023development] 0.6459$\pm$0.2219 0.3586$\pm$0.3677
EEGNet[lawhern2018eegnet] 0.6495$\pm$0.2063 0.3988$\pm$0.2848
deepConvNet[schirrmeister2017deep] 0.6854$\pm$0.1969 0.4539$\pm$0.3059
EEG2GAIT(w/o HGP) 0.7009$\pm$0.1822 0.4517$\pm$0.3337
EEG2GAIT 0.7126$\pm$0.1799 0.4835$\pm$0.304
Table 2: Prediction performance of different models on the on GED, indicated by mean $\pm$ standard deviation, averaged over the six joints. The highest mean value in each column is marked in bold, while the highest mean value in each row is marked with a superscript star (*). The lowest standard deviation in each column has been underlined.

### v2

@@ 3.4 Hierarchical GCN Pyramid
The HGP learns embedding $g^{i}$ for each subgraph $\hat{{X_{i}}}\in\mathbb{R}^{C\times F}$ to capture spatial features of the signal. A base graph encoder $\Phi_{g}(\cdot)$ is utilized to learn graph representations. In this paper, we use GCN [22] as the base graph encoder:
$\Phi_{g}(\mathbf{H}^{(l+1)},\mathbf{A})=\sigma\left(\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-\frac{1}{2}}\mathbf{H}^{(l)}\mathbf{W}^{(l)}+\mathbf{b}^{(l)}\right),$ (5)
@@
  • $\tilde{\mathbf{A}}$ represents the symmetrically normalized adjacency matrix $\hat{\mathbf{A}}$ computed during the forward pass from the learnable adjacency parameter $\mathbf{A}$;
  • $\tilde{\mathbf{D}}$ is the degree matrix used in the symmetric normalization process;
@@
The normalization process $\tilde{\mathbf{D}}^{-\frac{1}{2}}\tilde{\mathbf{A}}\tilde{\mathbf{D}}^{-\frac{1}{2}}$ is performed dynamically during each forward pass to ensure proper scaling of the learned adjacency matrix, where isolated nodes (degree 0) are handled by adding a mask to prevent division by zero.
To modulate the diversity of brain region connections for different cognitive processes, we propose to use multiple hierarchical GCNs, i.e., $\{\Phi^{0}_{g}(\cdot),\Phi^{1}_{g}(\cdot),\dots,\Phi^{i}_{g}(\cdot)\}$, with learnable adjacency matrices $\{A^{0},A^{1},\dots,A^{i}\}$. Each $A^{i}$ is initialized from spatially-informed priors but learns task-specific graph connections corresponding to a particular cognitive process during training. The hierarchical behavior is achieved by constructing $K$ parallel GCN branches with different depths $\{L^{0},L^{1},\dots,L^{K-1}\}$ where $L^{0}<L^{1}<\dots<L^{K-1}$, while maintaining consistent hyperparameters across all branches. For each $\Phi^{i}_{g}(\cdot)$, stacking different GCN layers allows learning different levels of node cluster similarity. Intuitively, for localized connections, such as electrodes within the same brain functional area, a deeper GCN can achieve consistent representations among these nodes. However, for global connections across different brain regions, a shallower GCN can aggregate information between these areas without oversmoothing. The deeper $\Phi^{i}_{g}(\cdot)$ becomes, the higher the level in the feature pyramid its output represents.
@@ 3.5 Global Spatial Learner
The GSL processes these combined features using a depth-wise grouped convolution [23] with a kernel size of $(C,1)$. This operation efficiently learns spatial filters specific to each temporal filter, enabling the extraction of frequency-specific spatial patterns. To further enhance the feature representations, batch normalization [19] is applied along the feature map dimension, followed by an exponential linear unit (ELU) [7] activation function. Dropout [44] with a probability of 0.5 is used to prevent overfitting when training on small datasets. Finally, an average pooling layer of size $(1,3)$ reduces the temporal resolution. As a result, the output size of the GSL is $(F,1,T//3)$.
@@ 3.6 Feature Fusion Layers
After the three convolutional blocks, the output passes through a dropout layer with a rate of 0.5 to mitigate overfitting. The final output of the Feature Fusion Layers has a size of $(200,1,T//81)$, ready for subsequent temporal modeling. This channel dimension (200) results from the 200 kernels used in the final convolutional block.
@@ 3.7 Global Temporal Learner
The Global Temporal Learner (GTL) captures temporal dynamics using a Multi-head Self-Attention (MSA) mechanism [47], which preserves the input-output shape consistency for efficient processing. Residual connections are employed to maintain the integrity of the input data and ensure stable gradient flow during training.
The input to the GTL is a tensor $X_{\text{spatial}}$ with shape $(200,T//81)$, representing the spatial-temporal features extracted from previous layers. To model temporal dependencies, $X_{\text{spatial}}$ is transformed into query ($Q_{i}$), key ($K_{i}$), and value ($V_{i}$) matrices for each attention head, where $Q_{i}$, $K_{i}$, and $V_{i}$ have dimensions $(T//81,d_{K})$. The self-attention mechanism for the $i$-th head is computed as:
$\text{head${}_{i}$}=\text{softmax}\left(\frac{Q_{i}\cdot K_{i}^{T}}{\sqrt{d_{K}}}\right)V_{i}$ (6)
@@
$X_{\text{attention}}=\text{linear}(\text{Concat}(\text{head}_{1},\ldots,\text{head}_{h}))$ (7)
Here, $h$ is the number of attention heads and $d_{K}$ is the output dimension of each attention head.
@@ 3.9.1 Time-Domain Loss
$L_{\text{time}}=\frac{1}{n}\sum_{i=1}^{n}\left(\hat{y}_{i}-y_{i}\right)^{2}$ (8)
@@
$L_{\text{time\_reward}}=L_{\text{time}}+\beta\cdot\log\left(1-e^{-L_{\text{time}}}+\epsilon\right)$ (9)
@@ 3.9.2 Frequency-Domain Loss
$L_{\text{freq}}=\frac{1}{n}\sum_{i=1}^{n}\left|\hat{Y}_{i}-Y_{i}\right|$ (10)
where $\hat{Y}(f)=\text{DFT}(\hat{y}(t))$ and $Y(f)=\text{DFT}(y(t))$ are the predicted and true signals in the frequency domain. Similar to the time-domain loss, a Reward Loss term is applied:
$L_{\text{freq\_reward}}=L_{\text{freq}}+\beta\cdot\log\left(1-e^{-L_{\text{freq}}}+\epsilon\right)$ (11)
@@ 3.9.3 HTSR Loss
$L_{\text{total}}=\alpha\cdot L_{\text{freq\_reward}}+(1-\alpha)\cdot L_{\text{time\_reward}}$ (12)
@@ 3.10.1 Gait-EEG Dataset
To investigate the neural mechanisms underlying natural gait, we collected a novel multimodal dataset, the Gait-EEG Dataset (GED), comprising synchronized recordings of brain activity and lower-limb joint kinematics during level-ground walking. The dataset includes recordings from 50 able-bodied participants (25 males, 25 females; aged 21 to 46, mean age 28.4, standard deviation 5.2), none of whom reported a history of neurological disorders or lower-limb pathologies. Participants engaged in two independent level-ground walking sessions, with every session comprising three identical walking blocks. Each block included approximately 40 trials, with each trial representing EEG signals and synchronized lower-limb joint angles as the participant walked straight for 7.7 meters. The experiment protocol for each block is shown in Fig. 2. Sessions were spaced at least three days apart.
Figure 2: Experiment protocol for one walking block.
Neural and Kinematic Data Acquisition
EEG signals were recorded using a 60-channel active EEG system (ActiCHamp Plus, Brain Products GmbH, Germany), along with 4 electrooculogram (EOG) channels to monitor ocular artifacts. Electrodes were positioned according to the extended 10–10 international system and referenced at FCz. The sampling rate was 1000 Hz. Participants wore an EEG cap and were assisted by a lightweight walking aid to reduce upper-body movement without interfering with natural gait. To further reduce motion-induced artifacts, electrode cables were secured, and participants wore tight-fitting caps and long compression pants provided by the experimenters.
Lower-limb joint angles at the hips, knees, and ankles were measured bilaterally using six twin-axis wireless goniometers (Biometrics Ltd, UK), also sampled at 1000 Hz. Sensor placement was verified through anatomical palpation at the start of each session and securely affixed using medical and kinesiology tape. The tight-fitting pants ensured both comfort during walking and minimal sensor displacement throughout the experiment.
Ethical Approval
This study was approved by the Institutional Review Board of Nanyang Technological University (IRB-2021-709), in accordance with applicable legislation, ethical guidelines, and safety regulations in Singapore. All participants provided written informed consent prior to participation.
Gait Cycle Segmentation
The gait cycles were segmented by detecting peaks in the left knee joint angle signal recorded by the goniometer. Each local maximum peak was considered as a boundary between consecutive gait cycles. The segment of joint angle data between two consecutive peaks thus defined one gait cycle.
@@ 3.10.2 Open-access Dataset
To further validate our proposed method, we also conducted experiments using an open-access dataset, Mobile Brain-body imaging(MoBI) dataset [15]. The MoBI dataset includes recordings from eight participants across three sessions each. For each session, participants walked at a constant speed on a treadmill for a total of 20 minutes. During the first 15 minutes, participants received real-time feedback of their gait via an on-screen avatar mirroring their actual movements. In the final 5 minutes, the avatar’s gait feedback was driven by decoded EEG signals.
4 Experiment
@@ 4.1 Data Preprocessing
4.1.1 Conventional Preprocessing
Prior to analysis, the EEG data from both datasets were subjected to identical preprocessing procedures. First, a minimum-phase band-pass filter (0.1–48 Hz) was applied, followed by re-referencing to the common average. The signals were then downsampled to 100 Hz. Finally, the electrooculography (EOG) channels were removed, resulting in EEG datasets with 59 channels for both the GED and MoBI datasets.
4.1.2 Local Laplacian Filtering
A local Laplacian filter was applied to enhance local activity within a specified area. It achieves this by computing the differences between the values of the target channel and the average of values from its neighboring channels. For each EEG channel, neighbors within a 30 mm radius were retained. This method effectively emphasizes local features while diminishing the interference from distant sources. The filtering process is expressed as:
@@ 4.1.2 Local Laplacian Filtering
where $V^{\prime}_{k}$ is the Laplacian-filtered value at channel $k$, $V_{k}$ is the original value, $N$ is the number of neighbors, and $V_{l}$ are the neighboring channel values.
4.1.3 ICA-based Artifact Removal (Control Experiment)
To further examine the potential influence of motion- or eye-related artifacts, we conducted a control experiment using Independent Component Analysis (ICA). All independent components obtained from ICA were subjected to the computation of five established metrics [29]: Spatial Average Difference, Spatial Eye Difference, Generic Discontinuities Spatial Feature, Maximum Epoch Variance, and Temporal Kurtosis.
A threshold for each metric was determined through maximum likelihood estimation. If any of the five metrics for a given component exceeded three times the threshold, that component was discarded. On average, 8.96 components were removed per session in the GED dataset and 9.54 components per session in the MoBI dataset. After artifact ICs were removed, the EOG channels were excluded, yielding a final EEG dataset of 59 channels for both datasets.
@@ 4.2.1 Gait-EEG Dataset
Each session comprises three blocks, with each block containing 40 trials. For data partitioning, we used all trials from the first two blocks and the first 20 trials of the third block as the training set. Trials 21 to 25 of the third block were designated as the validation set, while the remaining trials (26 to 40) of the third block were concatenated along the temporal dimension to construct the test set. As EEG2GAIT is designed as a session-specific model, we independently trained an EEG2GAIT model for each session using its respective training and validation sets, and evaluated it on the corresponding session’s test set.
@@ 4.2.2 MoBI Dataset
For each session, the first 15 minutes of walking data without closed-loop BCI control were allocated for model development. Within this portion, the initial 13.5 minutes were used for training the model, while the remaining 1.5 minutes (i.e., the last 10% of this segment) were reserved for validation. The final 5 minutes of walking data with closed-loop BCI control were kept as the independent test set.
@@ 4.3 Evaluation Metric
We evaluated the efficacy of EEG2GAIT by comparing the predicted angles of six joints with their actual recorded angles, using their correlation ($r$ value) and coefficient of determination ($R^{2}$ score) and the mean absolute error (MAE) as the evaluation metrics:
@@
$MAE=\frac{1}{n}\sum_{i=1}^{n}\left|y_{i}-\hat{y}_{i}\right|$ (16)
where $y$ denotes the actual joint angle, and $\hat{y}$ represents the predicted angle. The covariance between any two variables A and B is expressed as $cov(A,B)$, while $\sigma(A)$ signifies the standard deviation of A. Each sequence represents the data collected from a single trial, capturing the observations over a specific period. In this case, the sequence contains $n$ data points, which corresponds to $\frac{n}{100}$ seconds of data, given that the data was sampled at a frequency of 100 Hz. $\hat{y}_{i}$ refers to the value at the $i^{th}$ position. Lastly, $\bar{y}$ stands for the average of the actual joint angles. The $r$ value reflects the consistency in trend between the predicted and actual gait. The $R^{2}$ score reflects the second-order absolute error between the predicted and actual gait, providing insight into the predictive performance of the model. The MAE quantifies the average magnitude of the prediction error, offering a straightforward measure of how close the predictions are to the true values.
@@ 4.4 Implementation and Hyperparameter Settings
Table 1: Prediction performance of different models on the MoBI dataset, indicated by mean $\pm$ standard deviation, averaged over the six joints. The best mean value in each column is marked in bold, while the best mean value in each row is marked with a superscript star (*). For MAE, a lower value is better ($\downarrow$).
@@
$L_{time+freq}$ $L_{time\_reward}$ $L_{MSE}$
[…]
MAE results ($\downarrow$):
[…]
$L_{time+freq}$ $L_{time\_reward}$ $L_{MSE}$
[…]
5 Results and Analysis
Figure 3: Visualization of the actual and predicted six joint angles for the test set in participant 2’s second session. The goniometer-measured angles were segmented into gait cycles using peak detection, with each gait cycle interpolated to 400 time points. The blue solid line represents the actual mean angle, and the blue shaded area indicates the standard deviation of the measured angles across all gait cycles. The red dashed line denotes the predicted mean angle, while the red shaded area shows the standard deviation of the predictions.
In this section, we validate the performance of the proposed method on the Gait-EEG dataset and MoBI dataset [15] and compared with several state-of-the-art deep learning and machine learning algorithms in the brain-computer interface (BCI) domain, including ContraWR [52], FFCL [27], TSception [11], Temporal Convolutional Network (TCN) [18], ST-Transformer [42], EEGConformer [43], SPaRCNet [20], EEGNet [25], and deepConvNet [38]. An ablation study was conducted to demonstrate the contributions of the HGP module and HTSR loss. Saliency mapping across EEG channels identified specific channels with notably higher contributions to the decoding task, offering a spatial interpretation of the neural correlates involved in lower-limb motor decoding as well as validating the obtained features. This confirms that the model effectively leverages signals from motor-related areas. Additionally, we illustrate the overlay of decoding results across different models for the gait cycle, offering a more intuitive display of EEG2GAIT’s superior performance.
Fig. 3 illustrates the visualization of the predicted and actual six joint angles for the left knee during the test set of participant 2’s second session ($r$ = $0.978$, $R^{2}$ = $0.955$). To standardize the input for analysis and modeling, each gait cycle was temporally normalized by interpolating the angle data to a fixed length of 400 time points, enabling consistent comparison across different cycles and participants.
Table 2: Prediction performance of different models on GED, indicated by mean $\pm$ standard deviation, averaged over the six joints. The best mean value in each column is marked in bold, while the best mean value in each row is marked with a superscript star (*). For MAE, a lower value is better ($\downarrow$).
$r$ value results:
[…]
$L_{time+freq}$ $L_{time\_reward}$ $L_{MSE}$
[…]
$R^{2}$ score results:
[…]
$L_{time+freq}$ $L_{time\_reward}$ $L_{MSE}$
[…]
MAE results ($\downarrow$):
[…]
$L_{time+freq}$ $L_{time\_reward}$ $L_{MSE}$
[…]
5.1 Model Performance
The results of all metrics on the test sets are presented to evaluate the performance of EEG2GAIT in comparison to baseline methods on the MoBI and GED dataset. EEG2GAIT achieved an $r$ value of 0.959 ($R^{2}$ = 0.914) on the GED dataset with 50 subjects and an $r$ value of 0.779 ($R^{2}$ = 0.597) on the MoBI dataset with 8 subjects, outperforming the baseline methods. All baseline methods were trained according to the strategy described in Section 4.2. A summary of the results for EEG2GAIT and the baseline methods on the different datasets is provided in Table 1 and 2. Overall, EEG2GAIT demonstrated the best performance across all metrics among all methods, with a lower standard deviation than most methods. In addition, the decoding performance of EEG2GAIT for each individual joint across the three evaluation metrics ($r$, $R^{2}$, and MAE) is reported in Table 3.
Table 3: Joint-specific prediction results of EEG2GAIT model on GED across six lower-limb joints, indicated by mean $\pm$ standard deviation
Joint $r$ value $R^{2}$ score MAE
Left Hip 0.9677 $\pm$ 0.0302 0.9309 $\pm$ 0.0614 0.1761 $\pm$ 0.0741
Left Knee 0.9729 $\pm$ 0.0256 0.9437 $\pm$ 0.0522 0.1556 $\pm$ 0.0660
Left Ankle 0.9367 $\pm$ 0.0453 0.8669 $\pm$ 0.0902 0.2528 $\pm$ 0.0796
Right Hip 0.9674 $\pm$ 0.0348 0.9313 $\pm$ 0.0686 0.1744 $\pm$ 0.0639
Right Knee 0.9726 $\pm$ 0.0262 0.9391 $\pm$ 0.0768 0.1490 $\pm$ 0.0613
Right Ankle 0.9353 $\pm$ 0.0441 0.8730 $\pm$ 0.0801 0.2502 $\pm$ 0.0718